# Trade in Minutes! Rationality-Driven Agentic System for Quantitative Financial Trading

### v1

@@ 4.1 Implementation details
[…]
Grid M > 30m 0.16
ETF&PCA M > 7d -0.33
DDPG M > 12h 0.57
PatchTST M > 3d 0.67
FinMem M&N > 1d 0.41
TradingAgents M&N > 3d 0.58
TiMi (ours) M > 4h 0.91
4.2 Empirical results
@@ 4.3 Analytical study
Ablation study of the optimization stage. In Figure 5, the progression from the prototype $\mathcal{B}$ to the advanced $\mathcal{B}^{*}$ provides compelling support for the effectiveness of our optimization stage. Specifically, $\mathcal{B}$ persistently underperforms, hovering around break-even with final returns of merely 2%, while $\mathcal{B}^{*}$ achieves consistent growth culminating in over 20% CR. Meanwhile, the intermediate versions display valuable insights: $\mathcal{B}^{(1)}$ reaches promising peaks exceeding 35% CR in September but suffers from substantial drawdowns and eventual performance degradation, suggesting shallow parameter-level optimization is insufficient for sustained profitability. Conversely, $\mathcal{B}^{(3)}$ demonstrates more stable growth patterns, indicating the benefits of higher-level optimization. These findings align with the theoretical foundation introduced in Section 2.5, where the feedback reflection agent $\mathcal{A}_{\text{fr}}$ progressively refines trading strategies through constraint-based parameter solving and hierarchical interventions.
@@ 6 Conclusion
Reproducibility statement. To facilitate replication, we have provided further technical details in Appendix B, covering order management, exchange selection, transaction cost modeling, slippage handling, action latency control, and error handling mechanisms. Crucially, we commit to open-sourcing the implementation of TiMi for deployment and releasing the real transaction records from live trading (a preview is available in the Supplementary Material).
@@ Appendix B Details on Implementation Specifics for Deployment
We provide further implementation details of the TiMi system, covering order specifics, exchange selection, risk control mechanisms, transaction cost modeling, and action latency control. These technical details demonstrate the practical considerations necessary for deploying the system in live trading environments.
@@
Appendix C Ablation Study of Agent Components
@@
  • Strategy adaptation agent $\mathcal{A}_{\text{sa}}$: In Table 4, we provide the ablation results of a TiMi variant that bypasses pair-specific customization, simulated in the altcoin future market. The key insight is that while $\mathcal{A}_{\text{sa}}$ provides improvement in risk-adjusted returns, its most critical contribution (lower variance in returns across pairs) is enhancing robustness and performance consistency in a diverse market.
  • Feedback reflection agent $\mathcal{A}_{\text{fr}}$: The efficacy of $\mathcal{A}_{\text{fr}}$ is empirically validated by the ablation study in Figure 5. The prototype bot $\mathcal{B}$, lacking the optimization stage, stagnates near break-even performance. In contrast, the advanced bot $\mathcal{B}^{*}$, progressively refined by $\mathcal{A}_{\text{fr}}$, achieves consistent growth and a final return exceeding 20%. Crucially, $\mathcal{A}_{\text{fr}}$ achieves parameter solving by operating within the proposed hierarchical optimization scheme. As illustrated in Section 2.5 and visualized in Figure 2, the necessity of this scheme is confirmed by the unstable transient gains of bots after only shallow parameter-level tuning versus the sustained profitability of those refined at higher functional and strategic layers.
Table 4: Ablation study of the strategy adaptation agent $\mathcal{A}_{\text{sa}}$ on system performance. $\sigma_{\text{ARR}}$ represents the standard deviation of annualized returns across trading pairs, indicating cross-pair stability.
@@
TiMi  (w/ $\mathcal{A}_{\text{sa}}$) 13.7 0.86 32.8 11.0
TiMi (w/o $\mathcal{A}_{\text{sa}}$) 10.4 0.71 38.2 19.5

### v2

@@ 4.1 Implementation details
Agentic Implementation. We develop a hybrid communication protocol that combines XML-based message envelopes with JSON payloads to facilitate inter-agent data exchange. The XML layer encapsulates essential metadata (e.g., sender identity and contextual domain), while the JSON payload carries domain-specific content. Our agents operate within a deterministic environment equipped with system-level capabilities, including local file operations and verifiable API invocations. Crucially, we implement procedural posterior checks where TiMi programmatically verifies generated scripts and mathematical solutions within controlled sandboxes, capturing execution tracebacks to ensure computational outputs and parameter derivations satisfy predefined constraints before deployment.
@@
Table 1: Backtesting comparison on 2024 historical data across U.S. stock index futures and cryptocurrency markets. Optimal/suboptimal results are indicated by bold/underline, respectively.
Method U.S. Stock Index Futures Mainstream Coin Futures Altcoin Futures
[…]
Quantitative Methods
MACD 3.8 0.42 14.5 12.6 0.84 18.4 4.2 0.36 32.8
Momentum 8.0 0.73 16.8 15.7 0.93 21.4 -2.8 -0.40 38.3
Grid Trading -2.8 -0.35 10.2 -8.3 -0.72 22.6 9.4 0.72 18.7
Pairs Trading 2.3 0.40 7.4 -6.8 -0.67 16.3 -5.7 -0.46 14.5
ETF&PCA 4.6 0.49 13.9 7.2 0.62 15.8 6.3 0.69 17.2
TSMOM 10.5 0.81 19.7 18.4 1.15 17.2 1.5 0.05 41.6
OFI 2.1 0.34 11.8 8.9 0.66 19.3 7.8 0.64 23.9
ML&RL Methods
LSTM 3.2 0.36 15.3 9.4 0.88 17.8 -16.8 -0.89 25.4
DQN 8.3 0.69 18.6 9.7 0.83 20.5 -9.3 -0.80 38.7
DDPG 5.4 0.52 17.4 14.8 1.09 14.0 8.6 0.65 26.2
Autoformer 4.9 0.43 16.5 13.2 0.97 16.4 17.5 0.98 24.8
PatchTST 6.7 0.58 18.1 12.5 0.90 17.6 11.2 0.82 23.5
LLM-based Agents
FinGPT 5.8 0.57 15.7 6.6 0.59 19.8 7.4 0.61 31.2
FinMem 5.2 0.54 14.8 11.3 0.96 17.4 -8.9 -0.60 19.6
TradingAgents 6.3 0.60 16.2 17.9 1.12 20.3 9.7 0.72 29.6
TiMi (ours) 8.9 0.84 10.5 16.5 1.25 12.1 23.7 1.27 26.0
[…]
Grid M > 30m 0.16
ETF&PCA M > 7d -0.33
DDPG M > 12h 0.57
PatchTST M > 3d 0.67
FinMem M&N > 1d 0.41
TradingAgents M&N > 3d 0.58
TiMi (ours) M > 4h 0.91
4.2 Empirical results
Backtesting comparison. We evaluate on 2024 historical data to establish baseline performance. As detailed in Table 1, trend-following methods (e.g., TSMOM) capitalize on ETF-driven trends in mainstream coin markets, while LLM-agents exploit news and potential posterior information. Crucially, TiMi harmonizes high profitability with rigorous risk control, achieving superior stability and risk-adjusted returns (notably SR 1.27 in Altcoins). This highlights the robustness of our system in high-volatility, reflexive assets where traditional momentum or pure semantic analysis struggle.
@@ 4.3 Analytical study
Table 4: Component-wise ablation studies simulated within the 2024 crytocurrency market.
Method Configuration ARR% $\uparrow$ SR $\uparrow$ MDD% $\downarrow$ Live Deployment
TiMi full system 20.9 1.23 15.3 stable
$\mathcal{A}_{\text{fr}}^{\dagger}$ variant parameter-only optimization 12.5 0.92 16.3 logic inconsist
$\mathcal{A}_{\text{fr}}^{\ddagger}$ variant semantic-only reflection 1.1 0.05 25.1 stable
w/o $\mathcal{A}_{\text{sa}}$ unified strategy 15.2 0.95 28.4 stable
w/o $\mathcal{A}_{\text{fr}}$ prototype bot $\mathcal{B}$ 1.1 0.05 25.1 runtime unstable
w/o $\mathcal{A}_{\text{sa}}$ & $\mathcal{A}_{\text{fr}}$ minimal baseline -4.5 -0.21 34.2 runtime unstable
Ablation study of agents and optimization. Employing the macro analysis agent $\mathcal{A}_{\text{ma}}$ and bot evolution agent $\mathcal{A}_{\text{be}}$ as the operational backbone, we conduct granular component-wise ablation experiments simulated in the 2024 crytocurrency market to isolate the contributions of agent specialization (i.e., $\mathcal{A}_{\text{sa}}$ and $\mathcal{A}_{\text{fr}}$) and our optimization mechanism. As quantified in Table 4, removing $\mathcal{A}_{\text{sa}}$ nearly doubles Maximum Drawdown to 28.4%, underscoring its critical role in harmonizing diverse assets for consistent risk exposure (e.g., stable utility tokens vs. high-volatility meme coins). Regarding optimization, the semantic-only $\mathcal{A}_{\text{fr}}^{\ddagger}$ ensures stability via syntax resolution yet stagnates in profitability. Conversely, the parameter-only $\mathcal{A}_{\text{fr}}^{\dagger}$ yields theoretical gains but proves logically inconsistent for live deployment due to code-parameter dissonance, while the unoptimized baselines succumb to runtime instability. Consequently, only the full system sustains the policy-deployment-optimization chain, a synergy corroborated by the evolutionary trajectory in Figure 5; here, the unoptimized $\mathcal{B}$ stagnates near break-even and $\mathcal{B}^{(1)}$ suffers degradation despite transient 35% peaks, exposing the brittleness of shallow parameter tuning without structural adaptation, whereas the stabilized $\mathcal{B}^{(3)}$ culminates in $\mathcal{B}^{*}$ achieving consistent returns over 20%, validating the iterative efficacy of constraint-based solving and hierarchical interventions detailed in Section 2.5 and Figure 2.
@@ Appendix B Details on Implementation Specifics for Deployment
Reproducibility statement. To facilitate replication, we provide further implementation details of the TiMi system, covering order specifics, exchange selection, risk control mechanisms, transaction cost modeling, and action latency control. These technical details demonstrate the practical considerations necessary for deploying the system in live trading environments. Crucially, we will open-source the implementation of TiMi product for deployment and release the key corner case list (from both back simulation and live deployment).
@@
Appendix C Ablation Analysis of Agent Components
@@
  • Strategy adaptation agent $\mathcal{A}_{\text{sa}}$: In Table 6, we further provide the ablation results of $\mathcal{A}_{\text{sa}}$ simulated in the altcoin future market (2025). The key insight is that while $\mathcal{A}_{\text{sa}}$ provides improvement in risk-adjusted returns, its most critical contribution (lower variance in returns across pairs) is enhancing robustness and performance consistency in a diverse market.
  • Feedback reflection agent $\mathcal{A}_{\text{fr}}$: The efficacy of $\mathcal{A}_{\text{fr}}$ is empirically validated by the ablation study in Table 4 and Figure 5. The prototype bot $\mathcal{B}$, lacking the optimization stage, stagnates near break-even performance. In contrast, the advanced bot $\mathcal{B}^{*}$, progressively refined by $\mathcal{A}_{\text{fr}}$, achieves consistent growth and a final return exceeding 20%. Crucially, $\mathcal{A}_{\text{fr}}$ achieves parameter solving by operating within the proposed hierarchical optimization scheme. As illustrated in Section 2.5 and visualized in Figure 2, the necessity of this scheme is confirmed by the unstable transient gains of bots after only shallow parameter-level tuning versus the sustained profitability of those refined at higher functional and strategic layers.
Table 6: Ablation study of the strategy adaptation agent $\mathcal{A}_{\text{sa}}$ simulated in 2025 altcoin markets. $\sigma_{\text{ARR}}$ represents the standard deviation of annualized returns across pairs, indicating cross-pair stability.
@@
TiMi full system 13.7 0.86 32.8 11.0
w/o $\mathcal{A}_{\text{sa}}$ unified strategy 10.4 0.71 38.2 19.5